3.The following three charges are arranged as shown. Determine the net force acting on the charge on the far right (q3 = charge 3). q1 = 1.5 x 10-6C,q2 = - 2.3 x 10-6C,q3 = -3.5 x 10-3C.


4.The following charges are placed on the vertices of a right triangle. Charges are the same as that on problem 3.Find the net electrostatic force on charge 1.


1
Expert's answer
2022-03-10T18:01:32-0500

3.

"F=F_1+F_2=kq_3(\frac{q_1}{r_1^2}+\frac{q_2}{r_2^2})=126~kN,"

4.

"F=\sqrt{F_1^2+F_2^2}=kq_3\sqrt{\frac{q_1^2}{r_1^4}+\frac{q_2^2}{r_2^4}}=56~kN."